The following list shows the total expenditures in the private, public and foreign sectors in the United States in 2009 (in billions of dollars).


Household Consumption (C) = 10,001

Private Investment (I) = 1,590

Government Expenditures (G) = 2,914

Net Exports (X-M) = -386.

Calculate total US GDP for 2010. 




1
Expert's answer
2021-08-31T08:35:45-0400

Total US GDP for 2010 is:

"Y = C + I + G + (X - M) = 10,001 + 1,590 + 2,914 - 386 = 14,119"(billions of dollars) .